Can you please tell me if i did this right?
Solve for x:
7x - 3 = -5x - 18
add 5x to both sides which is 7x + 5x = 12x
and add 3 to both sides which is -18 +3 = -15
12x=-15
x= - 15/12 = 1 1/4

Answer by bucky(2189)

Yes. You did it correctly. Just to decrease your chances of making an error, you might want
to do it one step at a time as follows:
.
Given:
.
7x - 3 = -5x - 18
.
Add 5x to both sides to convert the equation to:
.
12x - 3 = -18
.
Add 3 to both sides to further reduce the equation to:
.
12x = -15
.
Solve for x by dividing both sides by 12 (which is the multiplier of x):
.
x = -15/12 = -1 1/4
.
But, what you did is just a shortened version of the above. You did nothing wrong and
your answer is correct. Good job! Keep up the good work!
.